OpenStudy (anonymous):

• What are the differences among expressions, equations, and functions? Provide examples of each.

OpenStudy (anonymous):

'2x-5y+5' is an expression whereas ' 2x-5y+5 = 0' is an equation. A function, on the other hand , is a completely different concept. An example could be f(x)=x^2

OpenStudy (anonymous):

An expression just shows you a set of operations with no equivalence. Such as a+b. An equation shows you sets of operations with an equivalence. Such as a+b=c. And a function take an independent variable (x) and maps it to a single value of y. Such as f(x)=x+2